# web_appendix

### v4

@@ 8 Conclusion
We propose new difference-in-difference (DID) estimators for continuous treatments. We assume that between pairs of consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units, the stayers, does not change. We propose a parallel trends assumption on the outcome evolution of switchers and stayers with the same baseline treatment. Under that assumption, two target parameters can be estimated. Our first target is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, referred to as the AS. Our second target is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, referred to as the WAS. Economically, the AS and WAS serve different purposes, so neither parameter dominates the other. On the other hand, when it comes to estimation, the WAS unambiguously dominates the AS. First, it can be estimated at the parametric rate even if units can experience an arbitrarily small treatment change. Second, under some conditions, its asymptotic variance is strictly lower than that of the AS estimator. Third, unlike the AS, it is amenable to doubly-robust estimation. In our application, we use US-state-level panel data to estimate the effect of gasoline taxes on gasoline consumption. The standard error of the WAS is almost three times smaller than that of the AS, and the two estimates are close.

### v5

@@ 7 Conclusion
We propose new, doubly-robust difference-in-difference (DID) estimators for treatments continuously distributed at every time period. We assume that between pairs of consecutive periods, the treatment of some units, the switchers, changes, while the treatment of other units, the stayers, does not change. We propose a parallel-trends assumption on the outcome evolution of switchers and stayers with the same baseline treatment. Under that assumption, two target parameters can be estimated. Our first target is the average slope of switchers’ period-two potential outcome function, from their period-one to their period-two treatment, referred to as the AS. Our second target is a weighted average of switchers’ slopes, where switchers receive a weight proportional to the absolute value of their treatment change, referred to as the WAS. The AS and WAS serve different purposes. When it comes to estimation, the WAS has two advantages. First, it can be estimated at the parametric rate even if units can experience an arbitrarily small treatment change. Second, under some conditions, the asymptotic variance of the WAS estimator is strictly lower than that of the AS. In our application, we use US-state-level panel data to estimate the effect of gasoline taxes on gasoline consumption. The standard error of the WAS is almost three times smaller than that of the AS, and the two estimates are close.
@@
Throughout, we assume that there are some stayers, namely units whose treatment does not change between consecutive periods. de Chaisemartin et al. (2024) discuss the extension of the results in this paper to applications without stayers.